OpenStudy (anonymous):

does anyone know how to find the derivative of e^2x/1+x?

OpenStudy (anonymous):

do you have a graphing calculator? if so, you can do it on there

OpenStudy (anonymous):

no i do not, my teacher wont let us use one

OpenStudy (anonymous):

T_T no i cant figure it out, sorry!

OpenStudy (anonymous):

its okay

OpenStudy (anonymous):

Use the quotient rule: \[\frac{d}{dx} (\frac{f(x)}{g(x)}) = \frac{g(x)*f'(x) - f(x)*g'(x)}{g(x)^{2}}\]

OpenStudy (anonymous):

Since you're given d/dx (\[\frac{e^{2x}}{1+x}\] \[f(x) = e^{2x} \ \ f'(x) = 2e^{2x} \ \ g(x) = 1+x \ \ g'(x) = 1\] Plug in and solve!